OpenStudy (anonymous):

Solve 2|5x − 2| + 4 = 18 for x.

OpenStudy (mathstudent55):

First get rid of the +4 from the left side but subtraction. Then get rid of the 2 to the left of the absolute value by division. Then you'll have an equation which is the absolute value of 5x - 2 = number. The next step is to separate it into two equations: Absolute val of 5x - 2 = some number OR abs value of 5x -2 = the negative of that number. Then solve both equations. the answer should be in the form x = __ OR x = __